### v1

# Combinatorial interpretation of a truncated identity of Gauss

Abstract. In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series. In particular, some results on a truncated identity of Guass have been proved. In this article, we provide a new combinatorial interpretation of the truncated identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
Keywords: truncated identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
AMS Classifications: 05A17, 11P84
1 Introduction
A partition $\pi$ of a positive integer $n$ is a finite non-increasing sequence of positive integers $\pi=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ such that $\pi_{1}+\pi_{2}+\cdots+\pi_{m}=n$. The empty sequence forms the only partition of zero. The $\pi_{i}$ are called the parts of $\pi$. Let $p(n)$ denote the number of partitions of $n$. The generating function of $p(n)$ was given by Euler
$\sum_{n=0}^{\infty}p(n)q^{n}=\frac{1}{(q;q)_{\infty}}.$
Here and in the sequel, we assume that $|q|<1$ and employ the standard notation:
$(a;q)_{\infty}=\prod_{i\geq 0}(1-aq^{i}),$
$(a;q)_{n}=\frac{(a;q)_{\infty}}{(aq^{n};q)_{\infty}},$
and
${M\brack N}=\left\{\begin{array}[]{ll}\frac{(q;q)_{M}}{(q;q)_{N}(q;q)_{M-N}},&\text{if }M\geq N\geq 0,\\ 0,&\text{otherwise.}\end{array}\right$.
One of the well-known theorems in the partition theory is Euler’s pentagonal number theorem:
$(q;q)_{\infty}=\sum_{n=0}^{\infty}(-1)^{n}q^{n(3n+1)/2}(1-q^{2n+1}).$ (1.1)
In 2012, Andrews and Merca [2] considered a truncated version of (1.1) and obtained that for $k\geq 1$,
$\frac{1}{(q;q)_{\infty}}\sum_{j=0}^{k-1}(-1)^{j}q^{j(3j+1)/2}(1-q^{2j+1})=1+(-1)^{k-1}\sum_{n=1}^{\infty}\frac{q^{k(k-1)/2+(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k-1}}$,
from which they deduced the following theorem for the partition function $p(n)$.
Theorem $1.$
[2, Theorem 1.1] For $n,k\geq 1$,
$(-1)^{k-1}\sum_{j=0}^{k-1}(-1)^{j}\left(p(n-j(3j+1)/2)-p(n-j(3j+5)/2-1)\right)=M_{k}(n),$
where $M_{k}(n)$ is the number of partitions of $n$ in which $k$ is the least integer that is not a part and there are more parts >k>k than there are <k<k.
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Guass [1, p. 23]:
$1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}=\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}.$ (1.2)
Inspired by Andrews and Merca’s work, Guo and Zeng [9] established the truncated version of (1.2) in 2013. They proved that for $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k}\sum_{n=k+1}^{\infty}\frac{(-q;q)_{k}(-1;q)_{n-k}q^{(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k}}$. (1.3)
An overpartition, introduced by Corteel and Lovejoy [7], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of 44.
$(4),(\overline{4}),(3,1),(\overline{3},1),(3,\overline{1}),(\overline{3},\overline{1}),(2,2),(2,\overline{2}),$
$(2,1,1),(\overline{2},1,1),(2,1,\overline{1}),(\overline{2},1,\overline{1}),(1,1,1,1),(1,1,1,\overline{1}).$
We impose the following order on the parts of an overpartition.
$\overline{1}<1<\overline{2}<2<\cdots.$ (1.4)
Let $\overline{p}(n)$ be the number of overpartitions of $n$. The generating function of $\overline{p}(n)$ is
$\sum_{n=0}^{\infty}\overline{p}(n)q^{n}=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}.$
The following corollary is an immediate consequence of (1.3).
Corollary 1.2.
[9, Corollary 1.2] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)\geq 0,$ (1.5)
with strict inequality if $n\geq(k+1)^{2}$. For example,
$\overline{p}(n)-2\overline{p}(n-1)\leq 0,\text{ with strict inequality if }n\geq 4.$ (1.6)
A combinatorial proof of (1.6) was given by Guo and Zeng [9]. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
In 2018, Andrews and Merca [6] provided the following version of (1.3). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(1-q^{k+j+1})(q^{k+j+2};q)_{\infty}},$
from which they deduced the following combinatorial interpretation of the sum in (1.5). The combinatorial proof of the following theorem was given by Ballantine, Merca, Passary and Yee [5].
Theorem 1.3.
[6, Corollary 8] For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=\overline{M}_{k}(n),$
where $\overline{M}_{k}(n)$ is the number of overpartitions of $n$ in which the first part larger than $k$ appears at least $k+1$ times.
In this article, we will present a new combinatorial interpretation of the sum in (1.5) by using the minimal excludant non-overlined part of an overpartition. The minimal excludant of a partition was introduced by Grabner and Knopfmacher [8] under the name “smallest gap”. Recently, Andrews and Newman [3] undertook a combinatorial study of the minimal excludant of a partition. The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$. There is a generalization of the minimal excludant of a partition. In [4], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo AA that is not a part of $\pi$. In [10], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
Definition 1.4.
[10, Definition 4.1] Let $\overline{mes}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo AA that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
By definition, it is easy to get that for $n\geq 1$,
$op_{A,a}(n)+\overline{op}_{A,a}(n)=\overline{p}(n).$ (1.7)
For example, we have
$\pi$ $(4)$ $(\overline{4})$ $(3,1)$ $(\overline{3},1)$ $(3,\overline{1})$ $(\overline{3},\overline{1})$ $(2,2)$ $(2,\overline{2})$
$\overline{mes}_{2,1}(\pi)$ 11 11 55 33 11 11 11 11
$\pi$ $(2,1,1)$ $(\overline{2},1,1)$ $(2,1,\overline{1})$ $(\overline{2},1,\overline{1})$ $(1,1,1,1)$ $(1,1,1,\overline{1})$
$\overline{mes}_{2,1}(\pi)$ 33 33 33 33 33 33
Hence, $op_{2,1}(4)=7$ and $\overline{op}_{2,1}(4)=7$.
For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{2,1}(\pi)\geq 2k+1$ and $\overline{mes}_{2,1}(\pi)\equiv 2k+1\pmod{4}$. By definition, we have
$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$ (1.8)
We give the following combinatorial interpretation of the sum in (1.5).
Theorem 1.5.
For $n,k\geq 1$,
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1).$
The following truncated version of (1.2) immediately follows from Theorem 1.5.
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).$
This article is organized as follows. In Section 2, we give a combinatorial proof of (1.6). Then, we show Theorem 1.5 in Section 3.
2 Combinatorial proof of (1.6)
Clearly, in order to prove (1.6), it is equivalent to showing that for $n\geq 1$,
$\overline{p}(n)/2\leq\overline{p}(n-1),\text{ with strict inequality if }n\geq 4.$ (2.1)
To do this, we introduce three sets of overpartitions. Bear in mind that the parts in an overpartition are ordered as in (1.4).
For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa. This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than 11 in $\lambda$ is greater than or equal to two plus the number of parts 11 in $\lambda$.
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than 11 in $\lambda$ is less than two plus the number of parts 11 in $\lambda$.
To give a combinatorial proof of (2.1), it suffices to show the following theorem.
Theorem 2.1.
There exists a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$ and $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$.
Proof. For $n\geq 1$, let $\pi=(\pi_{1},\ldots,\pi_{m-1},\pi_{m})$ be an overpartition in $\mathcal{A}(n)$. By definition, we know that $\pi_{m}$ is a non-overlined part, which implies that $\overline{1}$ does not appear in $\pi$. Assume that $\pi_{m}=t$, we consider the following two cases.
Case 1: $t=1$. In this case, we set $\lambda=(\pi_{1},\ldots,\pi_{m-1})$.
Case 2: $t\geq 2$. In this case, we set
$\lambda=(\pi_{1},\ldots,\pi_{m-1},\underbrace{1,\ldots,1}_{(t-2)^{\prime}\text{s}},\overline{1}).$
In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n)$. Obviously, the process above is reversible. Now, we have built a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$.
It remains to show that $\mathcal{C}(n-1)$ is nonempty for $n\geq 4$. It can be checked that for $n\geq 4$,
$(\overline{2},\underbrace{1,\ldots,1}_{(n-4)^{\prime}\text{s}},\overline{1})$
is an overpartition in $\mathcal{C}(n-1)$. This completes the proof.
3 Proof of Theorem 1.5
The objective of this section is to give a proof of Theorem 1.5. To do this, we need the following lemma.
Lemma 3.1.
For $n\geq 1$ and $j\geq 1$,
$\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1),$
Proof. For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ to $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$, and vice versa. It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$.
On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$. This completes the proof.
In [10], Yang and Zhou obtained the following theorem.
Theorem 3.2.
[10, Theorem 4.2] For $n\geq 1$, we have
$op_{2,1}(n)=\overline{p}(n)/2.$
Combining with (1.7) and (1.8), we can get the following lemma. We also give a proof of the following lemma with the aid of Lemma 3.1.
Lemma $3.$
For $n\geq 1$,
$\overline{p}(n)=2op_{2,1}(n,1).$
Proof. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}\right)=1.$ (3.1)
Comparing the coefficients of $q^{n}$ on the both sides of (3.1), we have
$\overline{p}(n)+2\sum_{j=1}^{\infty}(-1)^{j}\overline{p}(n-j^{2})=0.$
Combining with Lemma 3.1, we get
$\overline{p}(n)=-2\sum_{j=1}^{\infty}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)=2op_{2,1}(n,1).$
This completes the proof.
Now, we are in a position to give a proof of Theorem 1.5.
Proof of Theorem 1.5. By Lemmas 3.1 and 3.3, we have
$\quad(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)$
$=(-1)^{k}\left(2op_{2,1}(n,1)+2\sum_{j=1}^{k}(-1)^{j}(op_{2,1}(n,j)+op_{2,1}(n,j+1))\right)$
$=(-1)^{k}(-1)^{k}2op_{2,1}(n,k+1)$
$=2op_{2,1}(n,k+1),$
and thus the proof is complete.

### v2

@@
-Combinatorial interpretation of a truncated identity of Gauss
+Combinatorial interpretations of truncated versions of a identity of Gauss
 Abstract.
 In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series.
-In particular, some results on a truncated identity of Guass have been proved.
-In this article, we provide a new combinatorial interpretation of the truncated identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
-Keywords: truncated identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
+In particular, some truncated versions of a identity of Gauss have been proved.
+In this article, we provide new combinatorial interpretations of the truncated versions of the identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
+Keywords: truncated versions, a identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
 AMS Classifications: 05A17, 11P84
@@
 from which they deduced the following theorem for the partition function $p(n)$.
-Theorem $1.$
+Theorem 1.1.
 [2, Theorem 1.1] For $n,k\geq 1$,
@@
 where $M_{k}(n)$ is the number of partitions of $n$ in which $k$ is the least integer that is not a part and there are more parts >k>k than there are <k<k.
-Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Guass [1, p. 23]:
+Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, p. 23]:
 $1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}=\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}.$ (1.2)
@@
 $\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k}\sum_{n=k+1}^{\infty}\frac{(-q;q)_{k}(-1;q)_{n-k}q^{(k+1)n}}{(q;q)_{n}}{{n-1}\brack{k}}$. (1.3)
-An overpartition, introduced by Corteel and Lovejoy [7], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of 44.
+An overpartition, introduced by Corteel and Lovejoy [8], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of $4$.
 $(4),(\overline{4}),(3,1),(\overline{3},1),(3,\overline{1}),(\overline{3},\overline{1}),(2,2),(2,\overline{2}),$
@@
 Corollary 1.2.
-[9, Corollary 1.2] For $n,k\geq 1$,
+[11, Corollary 1.2] For $n,k\geq 1$,
 $(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)\geq 0,$ (1.5)
@@
 $\overline{p}(n)-2\overline{p}(n-1)\leq 0,\text{ with strict inequality if }n\geq 4.$ (1.6)
-A combinatorial proof of (1.6) was given by Guo and Zeng [9]. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
-In 2018, Andrews and Merca [6] provided the following version of (1.3). For $k\geq 1$,
-$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(1-q^{k+j+1})(q^{k+j+2};q)_{\infty}},$
+A combinatorial proof of (1.6) was given by Guo and Zeng [11]. In 2018, Andrews and Merca [3] provided the following truncated version of (1.2). For $k\geq 1$,
+$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
 from which they deduced the following combinatorial interpretation of the sum in (1.5). The combinatorial proof of the following theorem was given by Ballantine, Merca, Passary and Yee [7].
 Theorem 1.3.
-[6, Corollary 8] For $n,k\geq 1$,
+[3, Corollary 8] For $n,k\geq 1$,
 $(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=\overline{M}_{k}(n),$
 where $\overline{M}_{k}(n)$ is the number of overpartitions of $n$ in which the first part larger than $k$ appears at least $k+1$ times.
-In this article, we will present a new combinatorial interpretation of the sum in (1.5) by using the minimal excludant non-overlined part of an overpartition.
 The minimal excludant of a partition was introduced by Grabner and Knopfmacher [10] under the name “smallest gap”.
@@
 The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$.
+In [6], Ballantine and Merca proved some inequalities involving the minimal excludant of a partition.
+They also posed the following conjecture.
+Conjecture 1.4.
+[6, Conjecture 1] For $n,k\geq 1$,
+$\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-j(3j-1))\geq 0,$
+with strict inequality if $n\geq(k+1)^{2}$.
+Conjecture 1.4 was settled by Yao in [19]. Furthermore, Yao showed that for $\ell\geq 1$,
+$\sum_{n=0}^{\infty}\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-\ell j(3j-1)/2)q^{n}=2\frac{(q^{\ell};q^{\ell})_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}\frac{q^{(k+2j+1)^{2}}(1-q^{2k+4j+3})}{(q;q^{2})_{\infty}}.$ (1.7)
+In [11], Guo and Zeng conjectured a stronger inequality than (1.5).
+Conjecture 1.5.
+[11, (6.4)] For $n,k\geq 1$,
+$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})\geq 0,$ (1.8)
+with strict inequality if $n\geq k^{2}$.
+Conjecture 1.5 was proved independently by Mao [13] and Yee [20] and then reconfirmed later by Wang and Yee [14]. In [17], Xia, Yee and Zhao proved the following inequality which implies (1.8).
+$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k(k+1))\geq 0.$
+In [12], Li obtained the following truncated version of (1.2). For $k\geq 1$,
+$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$ (1.9)
+Li [12] and Wang and Xiao [15, 16] presented the combinatorial interpretations of the sum in (1.8).
+2 Main results
+In this section, we will list the main results in article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
 There is a generalization of the minimal excludant of a partition.
-In [4], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo AA that is not a part of $\pi$.
+In [5], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo $A$ that is not a part of $\pi$.
 In [18], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
 Definition 2.1.
-[10, Definition 4.1] Let $\overline{mes}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo AA that is not a non-overlined part in the overpartition $\pi$.
+[18, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$.
 For $n\geq 1$, define $op_{A,a}(n)$ (resp.
-$\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{A,a}(\pi)$ is congruent to $a$ (resp.
+$\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{A,a}(\pi)$ is congruent to $a$ (resp.
 $a+A$) modulo $2A$.
@@
 $\pi$ $(4)$ $(\overline{4})$ $(3,1)$ $(\overline{3},1)$ $(3,\overline{1})$ $(\overline{3},\overline{1})$ $(2,2)$ $(2,\overline{2})$
-$\overline{mes}_{2,1}(\pi)$ 11 11 55 33 11 11 11 11
+$\overline{mex}_{2,1}(\pi)$ 11 11 55 33 11 11 11 11
 $\pi$ $(2,1,1)$ $(\overline{2},1,1)$ $(2,1,\overline{1})$ $(\overline{2},1,\overline{1})$ $(1,1,1,1)$ $(1,1,1,\overline{1})$
-$\overline{mes}_{2,1}(\pi)$ 33 33 33 33 33 33
+$\overline{mex}_{2,1}(\pi)$ 33 33 33 33 33 33
 Hence, $op_{2,1}(4)=7$ and $\overline{op}_{2,1}(4)=7$.
-For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{2,1}(\pi)\geq 2k+1$ and $\overline{mes}_{2,1}(\pi)\equiv 2k+1\pmod{4}$.
+For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{2,1}(\pi)\geq 2k+1$ and $\overline{mex}_{2,1}(\pi)\equiv 2k+1\pmod{4}$.
 By definition, we have
 $op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$ (2.2)
-We give the following combinatorial interpretation of the sum in (1.5).
+In this article, we will present new combinatorial interpretations of the sum in (1.5) and (1.8) by using the minimal excludant non-overlined part of an overpartition.
 Theorem 2.2.
 For $n,k\geq 1$,
-$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1).$
-The following truncated version of (1.2) immediately follows from Theorem 1.5.
+$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1),$ (2.3)
+and
+$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=op_{2,1}(n,k)-op_{2,1}(n,k+1).$
+For $k\geq 1$, the generating function of $op_{2,1}(n,k+1)$ is
+$\begin{split}\sum_{n=1}^{\infty}op_{2,1}(n,k+1)&=(-q;q)_{\infty}\sum_{j=0}^{\infty}\frac{q^{1+3+\cdots+(2k+4j+1)}(1-q^{2k+4j+3})}{(q;q)_{\infty}}\\ &=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).\end{split}$ (2.4)
+Then, the following truncated version of (1.2) immediately follows from (2.3) and (2.4).
+Corollary 2.3.
+For $k\geq 1$,
 $\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).$
-This article is organized as follows. In Section 2, we give a combinatorial proof of (1.6). Then, we show Theorem 1.5 in Section 3.
+Combining Theorems 1.3 and 2.2, we can get the following corollary.
+Corollary 2.4.
+For $n,k\geq 1$,
+$op_{2,1}(n,k+1)=\frac{\overline{M}_{k}(n)}{2},$ (2.5)
+and
+$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=\frac{\overline{M}_{k-1}(n)}{2}-\frac{\overline{M}_{k}(n)}{2}.$ (2.6)
+Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.4), (2.5), and Euler’s partition identity [9] (see also [1, (1.2.5)]):
+$\frac{1}{(q;q^{2})_{\infty}}=(-q;q)_{\infty}.$
+Appealing to (1.9) and (2.6), we can get following corollary.
+Corollary 2.5.
+For $k\geq 1$,
+$\sum_{n=1}^{\infty}\left(\overline{M}_{k-1}(n)-\overline{M}_{k}(n)\right)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
+In this article, we will present an analytic proof of Corollary 2.5.
+This article is organized as follows.
+We first give a combinatorial proof of (1.6) in Section 3.
+Then, we show Theorem 2.2 in Section 4.
+Finally, we give an analytic proof of Corollary 2.5 in Section 5.
 3 Combinatorial proof of (1.6)
@@
 To do this, we introduce three sets of overpartitions. Bear in mind that the parts in an overpartition are ordered as in (1.4).
-For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined.
+For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For example, we have
+$\mathcal{A}(4)=\{(4),(3,1),(\overline{3},1),(2,2),(2,1,1),(\overline{2},1,1),(1,1,1,1)\}.$
 For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa.
 This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
-For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than 11 in $\lambda$ is greater than or equal to two plus the number of parts 11 in $\lambda$.
-For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than 11 in $\lambda$ is less than two plus the number of parts 11 in $\lambda$.
+For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$.
+For example, we have
+$\mathcal{B}(3)=\{(3),(\overline{3}),(2,1),(\overline{2},1),(2,\overline{1}),(1,1,1),(1,1,\overline{1})\}.$
+For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$.
+For example, we have
+$\mathcal{C}(3)=\{(\overline{2},\overline{1})\}.$
 To give a combinatorial proof of (3.1), it suffices to show the following theorem.
@@
 $\lambda=(\pi_{1},\ldots,\pi_{m-1},\underbrace{1,\ldots,1}_{(t-2)^{\prime}\text{s}},\overline{1}).$
-In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n)$.
+In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n-1)$.
 Obviously, the process above is reversible.
@@
 Lemma 4.1.
-For $n\geq 1$ and $j\geq 1$,
+For $n,j\geq 1$,
 $\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1),$
 Proof.
-For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ to $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$, and vice versa.
-It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$.
-On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$.
+For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ into $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$, and vice versa.
+It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$.
+On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$.
 This completes the proof.
@@
 Theorem 4.2.
-[10, Theorem 4.2] For $n\geq 1$, we have
+[18, Theorem 4.2] For $n\geq 1$, we have
 $op_{2,1}(n)=\overline{p}(n)/2.$
 Combining with (2.1) and (2.2), we can get the following lemma. We also give a proof of the following lemma with the aid of Lemma 4.1.
-Lemma $3.$
+Lemma 4.3.
 For $n\geq 1$,
@@
 $\overline{p}(n)=-2\sum_{j=1}^{\infty}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)=2op_{2,1}(n,1).$
-This completes the proof.
+The proof is complete.
 Now, we are in a position to give a proof of Theorem 2.2.
-Proof of Theorem 1.5. By Lemmas 3.1 and 3.3, we have
+Proof of Theorem 2.2. It follows from Lemma 4.1 that
+$\overline{p}(n-k^{2})=op_{2,1}(n,k)+op_{2,1}(n,k+1).$
+So, we just need to show (2.3). By Lemmas 4.1 and 4.3, we have
 $\quad(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)$
@@
 $=(-1)^{k}(-1)^{k}2op_{2,1}(n,k+1)$
-$=2op_{2,1}(n,k+1),$
-and thus the proof is complete.
+$=2op_{2,1}(n,k+1).$
+We arrive at (2.3), and thus the proof is complete.
+5 Analytic proof of Corollary 2.5
+In this section, we aim to give an analytic proof of Corollary 2.5. Note that the generating function of $\overline{M}_{k}(n)$ is
+$\sum_{n=1}^{\infty}\overline{M}_{k}(n)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
+so it remains to show
+$\begin{split}&\quad\frac{(-q;q)_{k-1}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.1)
+Dividing both sides of (5.1) by $\frac{(-q;q)_{k-1}}{(q;q)_{k}}$, we find that it is equivalent to
+$\begin{split}&\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.2)
+Then, we proceed to show (5.2).
+Proof of (5.2). Note that
+$\quad(1+q^{k})q^{(k+1)(k+j+1)}$
+$=q^{(k+1)(k+j+1)}+q^{(k+1)(k+j+1)+k}$
+$=q^{k(k+j+1)}(q^{k+j+1}-1+1)+q^{k(k+j+2)}(q^{k+j+1}+1-1)$
+$=q^{k(k+j+1)}(1-q^{k})-q^{k(k+j+1)}(1-q^{k+j+1})+q^{k(k+j+2)}(1+q^{k+j+1}),$
+so we get
+$\quad(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
+$=\sum_{j=0}^{\infty}\left((1-q^{k})\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+2};q)_{\infty}}\right$.
+$\qquad\qquad\qquad\left.+\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\right)$
+$=(1-q^{k})\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
+$\qquad\qquad\qquad+\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
+Then, we have
+$\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
+$=(1-q^{k})\frac{q^{k^{2}}(-q^{k+1};q)_{\infty}}{(q^{k};q)_{\infty}}+\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
+$=\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
+$=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
+We arrive at (5.2). This completes the proof.

### v3

@@
 In this article, we provide new combinatorial interpretations of the truncated versions of the identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
-Keywords: truncated versions, a identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
+Keywords: truncated versions, combinatorial interpretations, a identity of Gauss, the minimal excludant integer, overpartitions
 AMS Classifications: 05A17, 11P84
@@
 Let $p(n)$ denote the number of partitions of $n$.
-The generating function of $p(n)$ was given by Euler
+The generating function of $p(n)$ was given by Euler.
 $\sum_{n=0}^{\infty}p(n)q^{n}=\frac{1}{(q;q)_{\infty}}.$
 Here and in the sequel, we assume that $|q|<1$ and employ the standard notation:
-$(a;q)_{\infty}=\prod_{i\geq 0}(1-aq^{i}),$
+$(a;q)_{\infty}=\prod_{i=0}^{\infty}(1-aq^{i}),$
 $(a;q)_{n}=\frac{(a;q)_{\infty}}{(aq^{n};q)_{\infty}},$
@@
 where $M_{k}(n)$ is the number of partitions of $n$ in which $k$ is the least integer that is not a part and there are more parts >k>k than there are <k<k.
-Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, p. 23]:
+Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, (2.2.12)]:
 $1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}=\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}.$ (1.2)
@@
 Corollary 1.2.
-[11, Corollary 1.2] For $n,k\geq 1$,
+[10, Corollary 1.2] For $n,k\geq 1$,
 $(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)\geq 0,$ (1.5)
@@
 Conjecture 1.5.
-[11, (6.4)] For $n,k\geq 1$,
+[10, (6.4)] For $n,k\geq 1$,
 $(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})\geq 0,$ (1.8)
@@
 In [11], Li obtained the following truncated version of (1.2). For $k\geq 1$,
-$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$ (1.9)
-Li [12] and Wang and Xiao [15, 16] presented the combinatorial interpretations of the sum in (1.8).
+$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$ (1.9)
+based on which Li [11] presented a combinatorial interpretation of the sum in (1.8).
+Theorem 1.6.
+[11, Theorem 1.2] For $n,k\geq 1$,
+$(-1)^{k-1}\sum_{j=-k}^{k-1}(-1)^{j}\overline{p}(n-j^{2})=\overline{N}_{k}(n),$
+where $\overline{N}_{k}(n)$ is the number of overpartitions of $n$ in which a part of size $k$ has to be overlined, the smallest part >k1>k-1 appears exactly $k$ times and it cannot be overlined.
+In [14, 15], Wang and Xiao also obtained (1.9) and gave a combinatorial interpretation of the sum in (1.8). Li [11] generalized the sums in (1.5) and (1.8).
+Theorem 1.7.
+[11, Corollary 5.3] For integers $m$ and $k$ with $m\leq k$,
+$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})\geq 0.$ (1.10)
 2 Main results
-In this section, we will list the main results in article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
+In this section, we will list the results obtained in this article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
 There is a generalization of the minimal excludant of a partition.
@@
 Definition 2.1.
-[18, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$.
+[17, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest positive integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$.
 For $n\geq 1$, define $op_{A,a}(n)$ (resp.
@@
 Hence, $op_{2,1}(4)=7$ and $\overline{op}_{2,1}(4)=7$.
+Yang and Zhou [17] obtained the following theorem.
+Theorem 2.2.
+[17, Theorem 4.2] For $n\geq 1$, we have
+$op_{2,1}(n)=\overline{p}(n)/{2}.$ (2.2)
 For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{2,1}(\pi)\geq 2k+1$ and $\overline{mex}_{2,1}(\pi)\equiv 2k+1\pmod{4}$.
 By definition, we have
-$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$ (2.2)
-In this article, we will present new combinatorial interpretations of the sum in (1.5) and (1.8) by using the minimal excludant non-overlined part of an overpartition.
+$op_{2,1}(n)=op_{2,1}(n,0)\text{ and }\overline{op}_{2,1}(n)=op_{2,1}(n,1).$
+Combining with (2.1) and (2.2), we can get the following theorem.
 Theorem 2.3.
+For $n\geq 1$,
+$op_{2,1}(n,1)=\overline{p}(n)/{2}.$
+In this article, we will give another proof of Theorem 2.3.
+Then, we will present a combinatorial interpretation of the sum in (1.10) from the point of view of the minimal excludant non-overlined part of an overpartition.
+Theorem 2.4.
+For integers $m$ and $k$ with $m\leq k$,
+$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk>0,$ (2.3)
+and
+$(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,a+1)+(-1)^{m+k}op_{2,1}(n,b+1)\text{ if }mk\leq 0,$ (2.4)
+where $a=\text{min}\{|m|,|k|\}$ and $b=\text{max}\{|m|,|k|\}$.
+Setting m=k<0m=-k<0 and $m=-k+1\leq 0$ in (2.4), we can get new combinatorial interpretations of the sums in (1.5) and (1.8) respectively.
+Corollary 2.5.
 For $n,k\geq 1$,
@@
 $\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).$
-Combining Theorems 1.3 and 2.2, we can get the following corollary.
+Combining Theorem 1.3, Theorem 1.6 and Corollary 2.5, we can get the following corollary.
 Corollary 2.7.
 For $n,k\geq 1$,
-$op_{2,1}(n,k+1)=\frac{\overline{M}_{k}(n)}{2},$ (2.5)
+$op_{2,1}(n,k+1)=\overline{M}_{k}(n)/{2},$ (2.7)
 and
-$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=\frac{\overline{M}_{k-1}(n)}{2}-\frac{\overline{M}_{k}(n)}{2}.$ (2.6)
-Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.4), (2.5), and Euler’s partition identity [9] (see also [1, (1.2.5)]):
+$op_{2,1}(n,k)-op_{2,1}(n,k+1)=\overline{N}_{k}(n).$
+Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.6), (2.7), and Euler’s partition identity [1, (1.2.5)]:
 $\frac{1}{(q;q^{2})_{\infty}}=(-q;q)_{\infty}.$
-Appealing to (1.9) and (2.6), we can get following corollary.
+Note that $\overline{M}_{0}(n)=\overline{p}(n)$ for $n\geq 1$, then by Theorem 2.3, we have
+$op_{2,1}(n,1)=\overline{M}_{0}(n)/{2}.$
+Combining with Corollary 2.7, we get the following result.
 Corollary 2.8.
@@
 $\sum_{n=1}^{\infty}\left(\overline{M}_{k-1}(n)-\overline{M}_{k}(n)\right)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
-In this article, we will present an analytic proof of Corollary 2.5.
 This article is organized as follows.
 We first give a combinatorial proof of (1.6) in Section 3.
-Then, we show Theorem 2.2 in Section 4.
+Then, we show Theorems 2.3 and 2.4 in Section 4.
 Finally, we give an analytic proof of Corollary 2.8 in Section 5.
@@
 This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
-For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$.
+For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part lager than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$.
 For example, we have
 $\mathcal{B}(3)=\{(3),(\overline{3}),(2,1),(\overline{2},1),(2,\overline{1}),(1,1,1),(1,1,\overline{1})\}.$
-For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$.
+For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part larger than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$.
 For example, we have
@@
 is an overpartition in $\mathcal{C}(n-1)$. This completes the proof.
-4 Proof of Theorem 2.2
-The objective of this section is to give a proof of Theorem 2.2. To do this, we need the following lemma.
+4 Proofs of Theorems 2.3 and 2.4
+The objective of this section is to give the proofs of Theorems 2.3 and 2.4. To do this, we need the following lemma.
 Lemma 4.1.
 For $n,j\geq 1$,
-$\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1),$
+$\overline{p}(n-j^{2})=op_{2,1}(n,j)+op_{2,1}(n,j+1).$
 Proof.
@@
 This completes the proof.
-In [18], Yang and Zhou obtained the following theorem.
-Theorem 4.2.
-[18, Theorem 4.2] For $n\geq 1$, we have
-$op_{2,1}(n)=\overline{p}(n)/2.$
-Combining with (2.1) and (2.2), we can get the following lemma. We also give a proof of the following lemma with the aid of Lemma 4.1.
-Lemma 4.3.
-For $n\geq 1$,
-$\overline{p}(n)=2op_{2,1}(n,1).$
-Proof. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
+Then, we give a proof of Theorem 2.3 with the aid of Lemma 4.1.
+Proof of Theorem 2.3. Divide both sides of (1.2) by $\frac{(q;q)_{\infty}}{(-q;q)_{\infty}}$, we can get
 $\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{\infty}(-1)^{j}q^{j^{2}}\right)=1.$ (4.1)
@@
 The proof is complete.
-Now, we are in a position to give a proof of Theorem 2.2.
-Proof of Theorem 2.2. It follows from Lemma 4.1 that
-$\overline{p}(n-k^{2})=op_{2,1}(n,k)+op_{2,1}(n,k+1).$
-So, we just need to show (2.3). By Lemmas 4.1 and 4.3, we have
-$\quad(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)$
-$=(-1)^{k}\left(2op_{2,1}(n,1)+2\sum_{j=1}^{k}(-1)^{j}(op_{2,1}(n,j)+op_{2,1}(n,j+1))\right)$
-$=(-1)^{k}(-1)^{k}2op_{2,1}(n,k+1)$
-$=2op_{2,1}(n,k+1).$
-We arrive at (2.3), and thus the proof is complete.
+We conclude this section with a proof of Theorem 2.4.
+Proof of Theorem 2.4. We first show (2.3). There are the following two cases.
+Case 1: $k\geq m\geq 1$. In this case, we have $\text{min}\{|m|,k\}=|m|=m$, $\text{min}\{|m|,|k|\}=|m|=m$ and $\text{max}\{|m|,|k|\}=|k|=k$. It follows from Lemma 4.1 that
+$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
+$=(-1)^{m}\sum_{j=m}^{k}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
+$=(-1)^{m}\left((-1)^{m}op_{2,1}(n,m)+(-1)^{k}op_{2,1}(n,k+1)\right)$
+$=op_{2,1}(n,m)+(-1)^{m+k}op_{2,1}(n,k+1),$
+which agrees with (2.3) for $k\geq m\geq 1$.
+Case 2: $m\leq k\leq-1$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=-k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Again by Lemma 4.1, we have
+$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})$
+$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\overline{p}(n-j^{2})$
+$=(-1)^{k}\sum_{j=-k}^{-m}(-1)^{j}\left(op_{2,1}(n,j)+op_{2,1}(n,j+1)\right)$
+$=(-1)^{k}\left((-1)^{-k}op_{2,1}(n,-k)+(-1)^{-m}op_{2,1}(n,-m+1)\right)$
+$=op_{2,1}(n,-k)+(-1)^{m+k}op_{2,1}(n,-m+1).$
+We arrive at (2.3) for $m\leq k\leq-1$.
+Now, we have completed the proof of (2.3). Then, we proceed to show (2.4). Under the condition that $m\leq k$ and $mk\leq 0$, we have $m\leq 0\leq k$. Clearly, we have
+$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})+\overline{p}(n)+\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2}).$ (4.2)
+In view of (2.3), we get
+$\sum_{j=m}^{-1}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{m+1}op_{2,1}(n,-m+1)\right),$ (4.3)
+and
+$\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})=-\left(op_{2,1}(n,1)+(-1)^{k+1}op_{2,1}(n,k+1)\right).$ (4.4)
+Substituting (4.3) and (4.4) into (4.2) and combining with Theorem 2.3, we have
+$\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=(-1)^{m}op_{2,1}(n,-m+1)+(-1)^{k}op_{2,1}(n,k+1).$ (4.5)
+Then, we consider the following two cases.
+Case 1: $k\geq-m\geq 0$. In this case, we have $\text{min}\{|m|,k\}=|m|=-m$, $\text{min}\{|m|,|k|\}=|m|=-m$ and $\text{max}\{|m|,|k|\}=|k|=k$. Appealing to (4.5), we get
+$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,-m+1)+(-1)^{m+k}op_{2,1}(n,k+1).$
+So, (2.4) is valid for $k\geq-m\geq 0$.
+Case 2: $m>k\geq 0$. In this case, we have $\text{min}\{|m|,k\}=k$, $\text{min}\{|m|,|k|\}=|k|=k$ and $\text{max}\{|m|,|k|\}=|m|=-m$. Using (4.5), we get
+$\quad(-1)^{\text{min}\{|m|,k\}}\sum_{j=m}^{k}(-1)^{j}\overline{p}(n-j^{2})=op_{2,1}(n,k+1)+(-1)^{m+k}op_{2,1}(n,-m+1),$
+and thus (2.4) is satisfied for $m>k\geq 0$. This completes the proof.
 5 Analytic proof of Corollary 2.8
-In this section, we aim to give an analytic proof of Corollary 2.5. Note that the generating function of $\overline{M}_{k}(n)$ is
+In this section, we aim to give an analytic proof of Corollary 2.8. Note that for $k\geq 0$, the generating function of $\overline{M}_{k}(n)$ is
 $\sum_{n=1}^{\infty}\overline{M}_{k}(n)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$